Sol Summary – January 26th

Sol: 5
Summary Title: Data acquisition
Author’s name: Andrew Wheeler
Mission Status: Nominal
Sol Activity Summary: Sol 5 dawned bright clear and cold, perfect weather for EVAs. It was also Australia Day back on Earth and an opportunity for some Aussie delights. For us it was vegemite on toast with a tropical fruit platter as that was all we had. (The toast was actually bread and the fruit platter was dehydrated bananas, mango, apple and blueberries and the blueberries were actually in pancakes). Morning tea saw the unveiling of the TimTams® and, for afternoon tea, lamingtons. The hab received some tender loving care on both levels with sweeping up and vacuuming away the evidence of our external activities. The EVA for the day saw a return, along Cow Dung Road past Cowboy Corner, to the location of the spectral recordings of the previous day to collect further data and test the robot rover’s mobility on additional different slopes and surfaces. The evening meal was burgers and lamingtons. A productive day by all.
Look Ahead Plan: Weather dependent as always, tomorrow will see an EVA return to the Kissing Camel Ridge for investigation of the geological boundary between that and the Hab Ridge. This will be followed by an EVA to the Galileo Road in search of access into the Candor Chasma for geological samples and spectral targets.
Anomalies in work: Nothing to report.
Weather: Clear crisp morning at 0 Celcius rising to low teens during the day. Some minor clouds forming over the mountains during the day. No precipitation. Humidity dropping from near 80% to below 50%. The wind strengthened from gentle during the day.
Crew Physical Status: nominal
EVA: EVA 9 with Steve, Clare, and Rob to the Tharsis Montes area north of Cowboy Corner for robot rover testing.
Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, EVA Request(s). No Astronomer report as yet.
Support Requested: Nothing to request.

Operations Report – January 26th

SOL: 5
Name of person filing report: Scott Dorrington
Non-nominal systems: None
Notes on non-nominal systems: None
Spirit rover used:
Hours: 242.9
Beginning charge: 100%
Ending charge: 38%
Currently charging: No
Opportunity rover used:
Hours: 157.1
Beginning charge: 100
Ending charge: 46%
Currently charging: No
Curiosity rover used: Not used
Perseverance rover used: Not used
General notes on rovers: None
Summary of Hab operations:
WATER USE: 22.88 gallons
Water (static tank): 357.22 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off): on
Toilet tank emptied: no
Summary of internet: Nominal during comms window
Summary of suits and radios:
Suit #8 is still stood down out of commission – nuts missing from charging port. Opened up back of suit to see if nuts are in the bottom of suit. Did not find. Tomorrow, will locate new nuts to secure the charging port.
Summary of GreenHab operations:
WATER USE: 6.5 gallons
Heater: On
Supplemental light: Manually turned off at 9:30am. Returned to automatic control.
Harvest: None
Summary of ScienceDome operations: Not used
Dual split: Nominal
Summary of RAM operations: Entered RAM to retrieve replacement toilet seat
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support:
Thank you for the new toilet seat.

Journalist Report – January 26th

Australia Day
Sol 5 began with another interesting convo over breakfast – this time about human lung capacity and blood oxygen saturation, and whether/how a human could feasibly survive atmospheric depressurisation on Mars – at least, long enough to reach and operate an airlock. Somehow this led into a genealogy discussion about family trees, our ancestors, and their connections to Australia. Very fitting, given that today was Australia Day.
For international readers: Australia Day is the country’s official national holiday, observed by all its states and territories, and is one of the biggest events in the Australian calendar. The date of 26 January marks the anniversary of the British First Fleet’s landing near modern-day Sydney in 1788. While this specific date means different things to different Australians and has become somewhat controversial, for most it means a day off to come together and celebrate in solidarity. It’s also a day around which many citizenship and award ceremonies take place.
To commemorate the occasion, today was light-on with EVAs. We only conducted one, to Cowboy Corner again – but it was a doozy, pitting humans against Steve’s rover to compare their spectral capture and traversal capabilities. Humans won (the battle at least, maybe not the war). The team also took the opportunity for a flag-planting ceremony with the crew’s mascots.
Meanwhile, back at the hab, Commandrew was busy making improvised ‘lamingtons’ while dressed in his patriotic red, green and blue (see Pascal Lee’s Flag of Mars). Somehow they turned out great. We also baked more bread – at one point having three different loaves on the go – and improvised some ‘pseudo-snaggas’ that turned out… not-so-great.
Highlights of the day: None of the crew have known convict ancestors, Clare’s Scottish (Murray) side once owned a castle (Blair Atholl) with its own private army, Andrew’s eligibility to wear four different tartans, Rob’s Boer War family heirloom.
Lowlights of the day: Clare’s other side is a family of doppelgangers and she literally has no imagination (that’s not a joke), bread addiction, what happens on EVA stays on EVA.

– Alexander Tobal, Crew 291

Astronomy Report – January 29th



Name: Rob Hunt
Crew: 291
Date: 29/01/2024
Solar Features Observed: Sun spots (imaged) and solar prominences
Images submitted with this report: several images related to maintenance/repair, and an unprocessed one of the first 1,000 captures of Sun spots.
Problems Encountered:
entered observatory and removed cover from scope,
photographed position of scope with N, S, E in shot,
set up laptop and personal belongings, and decluttered work space (see pics),
a broken-off small, white, plastic positioning tab from the drawers was found in the lowest drawer and left there, (see pic)
dome was successfully rolled around to approximate position for use,
both shutters were opened without any glitches but, contrary to Peter’s direction to not use the lower shutter, it was required to be fully open in order to see the Sun which was low in the sky,
took more photos of pre-power scope use,
Musk Observatory Quick Guide (MOQG) cover title and elsewhere uses font/word art that is difficult to read,
MOQG instructions on p10 point to eyepiece attachment steps on p18 but they’re on p19 in chapter Visual Observing,
MOQG instructions on p10 point to camera attachment steps on p19 but they’re on p21 in chapter Solar Imaging Setup,
MOQG chapter Processing the Image on p27 is not listed in Table of Contents (ToC),
MOQG ToC chapter name for end of mission closing procedure doesn’t quite match chapter title,
MOQG chapters on Alignment Problems, Tuners, Focus Problems are not listed in ToC,
MOQG ToC chapter titles for Image Processing – Autostakkert, Registax and Photoshop don’t match chapter title,
switched on scope at pier, removed lens cap, then followed these instructions (deviation from instructions are in bold):
Once it’s at the index position it should go to the date and time routine,
I pressed ENTER and scope went to SWITCH POSITION then asked for time.
time and date (Standard and Mountain Time) were successfully entered,
If ALIGN SUCCESS doesn’t appear, and it asks for a 2-star alignment, then scroll (#6 and #9) to LAST ALIGNMENT and click ENTER,
I scrolled to LAST ALIGNMENT, clicked ENTER
Once you have ALIGN SUCCESS and its READY Click MENU,
I got ALIGN SUCCESS and CGE PRO READY then clicked MENU,
I scrolled to UTILITIES and got HOME POSITION and pressed ENTER,
In the HOME POSITION menu scroll until you get to SET HOME POSITION. Click ENTER,
there is no SET HOME POSITION in the HOME POSITION menu. There is only GOTO and SET,
I scrolled to SET and pressed ENTER,
the screen blinked while reading SETTING HOME POSITION then returned to SET in HOME POSITION,
Then scroll to HIBERNATE and shut down. It should operate normally then, but follow the shutdown procedure on page 15,
there is no GOTO HOME POSITION in the HOME POSITION menu but there is GOTO, so I clicked GOTO,
the scope slewed back and forth a bit and screen read GOTO in HOME POSITION menu,
there is no HIBERNATE in the HOME POSITION menu,
HIBERNATE is in the UTILITIES menu so I went there and completed the close-down procedure per MOQG p15,
scope was facing north and on top of the mount when I switched it off at the pier.
at approx. 14:45 it was very noticeable that the Sun was becoming fuzzy q.v. low in the sky and light travelling through much more atmosphere,
on speeds #4 and #3 (possibly others) the hand controller Direction Keys, if held down for a second or so, caused the motor to jump to high speed,
the lack of tightening screws in the eyepiece tube is workable but not ideal,
the USB cable to the laptop disconnected easily but was pushed back in securely and camera re-initiated in SharpCap,
prominences and Sun spots were seen and 1,000 images taken of the latter,
an unsuccessful attempt was made to capture a flat exposure,
closed observatory and returned to HAB.

Sol Summary – January 28th

Sol: 7
Summary Title: A day of rest
Author’s name: Andrew Wheeler
Mission Status: Nominal
Sol Activity Summary: Today was scheduled to be a day of rest and personal time. However, it didn’t quite turn out that way. One crew member awoke to a severe migraine and another to a stomach upset and their wellbeing was our main concern. Nevertheless, washing and cleaning was able to be done and a very short EVA to return the solar logger it’s recording point was carried out in the early morning. The Mars jigsaw puzzle was completed and signed and an issue with the lower level gas heater was remedied.
Look Ahead Plan: We will return to robot rover testing on the ground around the hab in the morning and an EVA to the stream draining from Tharsis Montes in search of concretions in the afternoon. If the weather holds, the solar observatory will also be operated.
Anomalies in work: Lower level gas heater anomaly – see operations report.
Weather: Clear crisp morning at -5 Celsius rising to low teens during the day. High cirrus cloud later in the day. No precipitation. Humidity continuously dropping from near 60% to below 40%. The wind was faint to gentle during the day.
Crew Physical Status: Severe migraine headache as well as stomach upset – see HSO report.
EVA: EVA 12 with Andrew and Clare to return the solar logger to it’s recording position and an external inspection of the hab tunnels.
Reports to be filed: Sol Summary, Commander’s Report, Greenhab Report, Operations Report (including an HSO report), EVA report, EVA Request(s). There is no Astronomer report for today.
Support Requested: Nothing to request.

Operations Report – January 28th

SOL: 7
Name of person filing report: Scott Dorrington
Non-nominal systems: Hab downstairs wall heater
Notes on non-nominal systems:
At around 18:30pm, the crew upstairs heard a strange sound coming from downstairs (sounded like rushing water). We went downstairs to investigate and smelled gas. We saw that the wall-mounted heater (to the right of the engineering airlock) that the left side of the heating grill was operating normally (glowing red), but the right was not glowing and there was a visible flame behind. Commander Andrew immediately turned the thermostat to the off position as I radioed mission support to advise on the situation. The flame then went out and the rushing sound stopped. We opened the engineering airlock door to vent the gas from the room. No alarms were triggered. The gas heater thermostat is now set to “off” and confirm the pilot light is still burning.
Spirit rover used: Not used
Opportunity rover used: Not used
Curiosity rover used: Not used
Perseverance rover used: Not used
General notes on rovers: None
Summary of Hab operations: put notes here
WATER USE: 29.4 gallons
Water (static tank): 301.7 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off): on
Toilet tank emptied: no
Summary of internet: Nominal during comms window
Summary of suits and radios: Nominal. Suit #8 still stood down pending new nuts in charging port.
Summary of GreenHab operations: Heater observed to be working nominally (turning on at 75 F)
WATER USE: 7 gallons
Heater: On
Supplemental light: 10pm – 2am
Harvest: None.
Summary of ScienceDome operations: Not used
Dual split: None
Summary of RAM operations: Not used
Summary of any observatory issues: Not used
Summary of health and safety issues:
Crew Physical Status: HSO – 1 X crew member suffering from gastro and 1 X crew member suffering from migraine. Both members are staying hydrated and undergoing bed rest. HSO has advised wearing gloves in living area to reduce infection transmission plus using the same cups, plates and cutlery.
Questions, concerns and requests to Mission Support:
Please advise any follow up instructions regarding the downstairs hab gas heater. Should we keep it turned off?

GreenHab Report – January 28th

GreenHab Officer: Scott Dorrington
Environmental control: heater on automatic, supplemental lights automatic 10pm to 2am
Average temperatures:
09:17: 73.6 F, 22%
09:59: 77.7 F, 25%
11:05: 83.8 F, 21%
12:04: 92.8 F, 19%
13:03: 97.3 F, 15%
13:51: 100.8 F, 12%
14:03: 89.1 F, 13%
15:27: 100.8 F, 11%
17:39: 76.8 F, 27%
Hours of supplemental light: 4 hours overnight
Daily water usage for crops: 7 gallons
Daily water usage for research and/or other purposes: 0 gallons
Water in Blue Tank (200 gallon capacity): 124 gallons
Time(s) of watering for crops:
Changes to crops: Planted two additional rows of carrots in the carrot box
Today was the crew’s rest day, although the plants still needed tending. I visited the GreenHab in the morning around 9:15 and stayed for a while to check in on the heater. It was on when I entered at 73 F and turned off several minutes later when the temperature read 76 F. Heating system nominal!!! I started the plants with a nice morning misting. Unfortunately, cukes #1 and #3 were looking a little wilty. I suspect they may have gotten some direct hot air from the heater – they were placed on the metal shelves overnight. I made a note not to leave them there when the heater was on. They were placed down on the floor for the day in the shade of the mesh shade cloth (later moved to the wooden pallet with the tomatoes). I also pruned some dead leaves from tomato #4 and inspected the fruits on tomato #9 – they look nice and red and a little squishy to the touch. I think tomorrow might be harvest day for them. I gave the plants their first watering of the day with 3 gallons. Soundtrack of the day was Jackson Browne. Today, I joined the crew in completing the Mars puzzle on our rest day. I set regular alarms to check in hourly on the plants, recording the temperature and giving a mist on each visit. By 1pm, the temperature had reached 100 F, although I did not observe the anomalous heater-turning-on issues the crew observed the previous sol. Looks like there may have been an initial hiccup in the system, but is now working ok. I gave the plants another watering of 3 gallons in the afternoon, and did some rotations of the tomatoes, and moving other plants out of the sun. Alex caught a bee inside the hab, which was also relocated to the GreenHab (I have yet to see it). For the past few sols, I have been eying off the carrot box – only a few carrot heads appear to have germinated, and there is a lot of free space in the box. Luckily, I did bring a packet of carrot seeds with me. I dug two shallow trenches on the left side of the box, away from the existing carrots, and added some seeds – following the instructions to apply 4 seeds per inch during initial planting, and to spread them out later after germination (expected to be 8-12 days). I will make sure to keep the carrot box nice and moist. In the afternoon, temperatures began to cool, and the heater was turning on periodically keeping the temperature above 75 F.
Harvest: None
Support/supplies needed: None

EVA Report – January 28th

EVA # 12
Author: Clare Fletcher
Purpose of EVA: Deploy solar logger before sun gets too high. Solar logger had stopped working on 27-01-24 and was retrieved on EVA #10.
Start time: 0800
End time: 0815
Narrative: Prior to the beginning of EVA #12, Alex reported having a migraine and was unable to join this EVA. Given I (Clare) was awake, I volunteered to help on this EVA to deploy the solar logger. Andrew and I made good time getting ready and left on time. We deployed the solar logger outside the Science Dome and turned it on. After this, we inspected the tunnels quickly as we intend to hold a practice emergency depressurization EVA during the week (EVA request yet to come for that one).
Destination: Outside Science Dome. 518150, 4250950
Coordinates (use UTM WGS 84): 518150, 4250950
Participants: Andrew Wheeler, Clare Fletcher (replacing Alex Tobal)
Road(s) and routes per MDRS Map: Exit Hab via main airlock and walk to outside the Science Dome. Within 50m of the Hab walking distance.
Mode of travel: Walking only.

Copyright © The Mars Society. All rights reserved. | Main Site